# Ecologically rational meta-learned inference explains human category learning

### v1

@@ 3.1 Prompting LLMs to generate ecologically valid category learning tasks
Synthesize feature names and category labels
I am a psychologist who wants to run a category learning experiment. In a category learning experiment, there are many different three-dimensional stimuli, each of which belongs to one of two possible real-world categories. Please generate names for three stimulus feature dimensions and two corresponding categories for 250 different category learning experiments:
@@
Generate category learning tasks
I am a psychologist who wants to run a category learning experiment. For a category learning experiment, I need a list of stimuli and their category labels. Each stimulus is characterized by three distinct features: sodium, fat, and protein. These features can take only numerical values. The category label can take the values healthy or unhealthy and should be predictable from the feature values of the stimulus. Please generate a list of 100 stimuli with their feature values and their corresponding category labels using the following template for each row: – feature value 1, feature value 2, feature value 3, - category label
@@
Figure 1: LLM generates ecologically valid category learning tasks: Mean task performance of the logistic regression model measured over trials (first column). Histogram of Pearson’s correlation coefficients computed between pairs of features (second column). Histogram of Gini coefficients computed over the logistic regression weights (third column). Linearity of the category learning task (fourth column) computed for (a) 28 different real-world binary classification tasks from the OpenML-CC18 benchmarking suite (b) ecologically valid category learning tasks generated from Claude-v2 and (c) synthetic category learning tasks derived using the Bayesian logistic regression prior.
@@ 3.2 Ecologically rational meta-learned inference (ERMI)
We refer to the class of models derived by training on ecologically valid (i.e., LLM-generated) category learning task as ecologically rational meta-learned inference (ERMI). If trained on synthetically-generated category learning tasks sampled from a Bayesian logistic regression prior, we refer to the models as meta-learned inference (MI; Binz et al., 2022a). Finally, using the terminology of Müller et al. (2022), we refer to the models as prior-data fitted networks (PFN) when tasks are sampled from a Bayesian neural network prior. For details on how these tasks are generated, see Appendix B.1.
@@ 4 LLM-generated category learning tasks are ecologically valid
We downsampled all tasks in the OpenML-CC18 benchmark to four feature dimensions and included only binary classification tasks without any missing features in our analysis – amounting to 28 tasks. In addition, we also contrasted LLM-generated tasks to a collection of synthetically-generated category learning tasks with a linear decision boundary (corresponding to those used to train MI). We analyzed these collections of tasks in terms of their learning curves, input correlations, sparsity, and linearity – details for which can be found in Appendix C.
@@ 5 ERMI shows human-like learning difficulties
We simulated ERMI and MI on the Shepard et al. (1961) study. Figure 2 (a) to (c) shows their learning curves alongside those of humans for the six difficulty levels. It can be seen that the learning curves of ERMI are difficulty-dependent and are in terms of mean-squared error (MSE) more similar to humans ($\mathrm{MSE}=0.03$) than MI ($\mathrm{MSE}=0.26$). Notably, ERMI, like humans, finds the type 1 task easier than the type 6 task and shows a similar clustering of learning curves for type 2 to type 5 tasks. However, we also find that ERMI learns much faster than people: its performance plateaus after four blocks while humans continue learning until the end of the experiment for most types. Even though MI performs tasks of different difficulty levels to varying degrees of success, its learning curves do not match those of humans. For example, unlike humans, MI finds type 2 tasks as difficult as type 6 tasks and type 4 tasks easier than type 3 tasks.
@@ 9 Discussion
Ecological rationality has a long history in cognitive science (Brunswik, 1955; Simon, 1990; Todd & Gigerenzer, 2012). However, it has been notoriously difficult to build models that are optimally adapted to the problems that people encounter in their everyday environment. From a technical perspective, the framework of meta-learning offers a solution to this problem. Yet, it has thus far only been applied to artificially-generated environments (Kumar et al., 2020; Binz & Schulz, 2022b; Binz et al., 2022a; Lake & Baroni, 2023; Jagadish et al., 2023). The main obstacle up to now was that the number of available real-world data sets was insufficient for meta-learning. We have shown that one can overcome this obstacle by prompting LLMs to generate a large collection of ecologically valid category learning tasks. We have then used meta-learning to obtain models that are optimally adapted to them, leading to a class of ecologically rational models that we call ERMI.
@@
However, at the same time, there were also facets of human category learning not captured by ERMI. In particular, we found that ERMI generally learned faster than people. In the third experiment, for instance, ERMI already displayed the generalization patterns shown in Figure 4 after two blocks, while people required 16 blocks or more (Johansen & Palmeri, 2002). We believe that this gap can (at least partially) be closed by incorporating limited computational resources (Binz et al., 2022a; Jagadish et al., 2023) or other architectural constraints (Achterberg et al., 2023).
To conclude, we have shown that LLMs can generate ecologically valid category learning tasks that can be used for meta-learning. With these models at hand, we then demonstrated that one can explain human category learning to a large extent using the principle of ecological rationality. Furthermore, the priors acquired by ERMI are rich enough that it achieves state-of-the-art performance on a real-world classification benchmark. In future work, we plan to scale up ERMI’s architecture, train it on classification tasks with a flexible number of features, increase the maximum number of data points, and allow for more than two classes.
@@ B.2 Data pre-processing, model architecture, and training
We filter out all tasks with more than two unique category labels and then binarize the category labels which are originally strings to make them consistent across tasks. We also normalized each feature independently using a min-max normalization scheme such that values taken by any feature lie always between zero and one. Both the task features and data points were shuffled while generating tasks. Note that the tasks generated by LLMs are typically of different lengths. Whenever the sampled tasks are of variable lengths, they are padded with zeros to match the length of the longest task sample within the batch. We additionally also sampled LLM-generated data points with replacement to match the length of the experimental task used in the Devraj et al. (2021) and Johansen & Palmeri (2002) studies. We resorted to this strategy as the LLM-generated tasks had a maximum of about 200 data points per task and by resampling, we can evaluate the model on experiments with larger horizons without any drop in performance. The batch size was set to 64 for three- and four-dimensional stimuli and to 32 for six-dimensional stimuli and it operated under a maximum steps regime of 400, 300, and 650 for three, four, and six-dimensional tasks respectively.
@@
The task features were mapped onto a 64-dimensional embedding space and positional encoded using sine and cosine functions of different frequencies as in Vaswani et al. (2017). A casual attention mask was then generated for the inputs such that the model makes conditional predictions on all preceding data points. The inputs along the attention mask are then passed to the transformer decoder model which had six layers, a model dimension of 64, 256 hidden units in the feed-forward network, and eight attention heads. The output of the transformer was then passed through a linear readout and sigmoid function to generate probability estimates for category 1. In practice, inference for all time steps is performed in parallel by passing a causal attention mask to the TransformerDecoder module in PyTorch (Paszke et al., 2019). We used binary cross-entropy (BCE) loss for a given batch of inputs and updated the model parameters using the ADAM optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-4}$. We trained all our models for a total of 500 000500000500\,000500 000 episodes.
@@ H.3 Additional observations and results
A better fit of the prototype model to the MI and PFN in the latter half of the experiment (as shown in Figure 3 and 8) suggests that, as observed in Shepard’s task, they are not able to learn exceptions like the prototype model as mentioned before.

### v2

# Human-like Category Learning by Injecting Ecological Priors from Large Language Models into Neural Networks

@@ 3.1 Prompting LLMs to generate ecologically valid category learning tasks
Synthesize feature names and category labels I am a psychologist who wants to run a category learning experiment. In a category learning experiment, there are many different three-dimensional stimuli, each of which belongs to one of two possible real-world categories. Please generate names for three stimulus feature dimensions and two corresponding categories for 250 different category learning experiments:
@@
Generate category learning tasks I am a psychologist who wants to run a category learning experiment. For a category learning experiment, I need a list of stimuli and their category labels. Each stimulus is characterized by three distinct features: sodium, fat, and protein. These features can take only numerical values. The category label can take the values healthy or unhealthy and should be predictable from the feature values of the stimulus. Please generate a list of 100 stimuli with their feature values and their corresponding category labels using the following template for each row: – feature value 1, feature value 2, feature value 3, - category label
@@
Figure 1: LLM generates ecologically valid category learning tasks: Mean task performance of the logistic regression model measured over trials (first column). Histogram of Pearson’s correlation coefficients computed between pairs of features (second column). Histogram of Gini coefficients computed over the logistic regression weights (third column). Linearity of the category learning task (fourth column) computed for (a) 28 different real-world binary classification tasks from the OpenML-CC18 benchmarking suite (b) ecologically valid category learning tasks generated from Claude-v2 and (c) synthetic category learning tasks derived using the Bayesian logistic regression prior that were used to train the meta-learned inference (MI) model (d) synthetic category learning tasks with nonlinear decision boundary derived using the Bayesian neural network prior that were used to train prior-fitted networks (PFN) model.
@@ 3.2 Ecologically rational meta-learned inference (ERMI)
We refer to the class of models derived by training on ecologically valid (i.e., LLM-generated) category learning task as ecologically rational meta-learned inference (ERMI). If trained on synthetically-generated category learning tasks sampled from a Bayesian logistic regression prior, we refer to the models as meta-learned inference (MI; Binz et al., 2022a). We chose MI as a baseline for two reasons. First, linear models have a long history as models of human learning (Karelaia & Hogarth, 2008; Lucas et al., 2015) and also category learning more specifically (Speekenbrink et al., 2008, 2010). Furthermore, the model has been previously used in a setting known as multiple cue probability learning (Binz et al., 2022a). While not exactly the same, this setting shares many similarities to category learning. Finally, using the terminology of Müller et al. (2022), we refer to the models as prior-data fitted networks (PFN) when tasks are sampled from a Bayesian neural network prior. For details on how these tasks are generated, see Appendix B.1.
@@ 4 LLM-generated category learning tasks are ecologically valid
We downsampled all tasks in the OpenML-CC18 benchmark to four feature dimensions and included only binary classification tasks without any missing features in our analysis – amounting to 28 tasks. In addition, we also contrasted LLM-generated tasks to a collection of synthetically-generated category learning tasks with a linear decision boundary (corresponding to those used to train MI; refer to Appendix B.1). We analyzed these collections of tasks in terms of their learning curves, input correlations, sparsity, and linearity – details for which can be found in Appendix C.
@@ 5 ERMI shows human-like learning difficulties
We simulated ERMI and MI on the Shepard et al. (1961) study. Figure 2 (a) to (c) shows their learning curves alongside those of humans for the six difficulty levels. It can be seen that the learning curves of ERMI are difficulty-dependent and are in terms of mean-squared error (MSE) more similar to humans ($\mathrm{MSE}=0.03$) than MI ($\mathrm{MSE}=0.26$). Notably, ERMI, like humans, finds the type 1 task easier than the type 6 task and shows a similar clustering of learning curves for type 2 to type 5 tasks (see Table 5 for details). However, we also find that ERMI learns much faster than people: its performance plateaus after four blocks while humans continue learning until the end of the experiment for most types. Even though MI performs tasks of different difficulty levels to varying degrees of success, its learning curves do not match those of humans. For example, unlike humans, MI finds type 2 tasks as difficult as type 6 tasks and type 4 tasks easier than type 3 tasks.
@@
Additionally, we simulated Claude-v2 directly on the Badham et al. (2017) study (see Appendix G for details). We first observed that Claude-v2 can perform the task optimally quite rapidly, with a mean error reaching zero for most difficulty levels already after three blocks. However, we also observed that the ordering of difficulty levels in Claude-v2 does not match humans, with Claude-v2 finding type 2 tasks more difficult than type 3, 4, and 5 tasks as shown in Figure 10. Furthermore, we found that the LLM’s (MSE=$0.18$) match to human error rates was worse than that of ERMI (MSE=$0.05$), see Table 6 for details. We also fitted the Claude-v2 predictions to human choices from the Badham et al. (2017) study and compared its performance to ERMI using the Bayesian information criterion (BIC; lower is better). We found that ERMI (38722.13) still offers a better fit to humans than Claude-v2 (39816.90). Taken together, these results suggest that an LLM trained on the entire internet cannot explain human category learning as well as an ecologically rational meta-learning inference model.
@@ 6 ERMI becomes more exemplar-based with learning
We additionally simulated Claude-v2 directly on the Devraj et al. (2021) study. We found that Claude-v2 is completely exemplar-based after the second block, which significantly differs from the strategy used by humans and ERMI (see Figure 12). We also found that ERMI (41207.20) offers a better fit to human choices than Claude-v2 (42401.42) in terms of BIC. These results once again confirm that while Claude-v2 can generate ecologically valid data, it cannot bring to bear its rich set of ecological priors to act in human-like ways.
@@ 8 ERMI achieves state-of-the-art performance on machine learning benchmarks
We additionally adapted a Bayesian analysis from (Stephan et al., 2009) to compute the probability that a model offers the best performance (within a set of alternative models) most frequently across data sets. This measurement is also known as the exceedance probability (EXP) in the statistics literature and is reported in the last row of Table 1. We found that ERMI is the best model according to this metric, with an exceedance probability of 0.66. TabPFN comes second with an exceedance probability of 0.24.
@@ 9 Discussion
Ecological rationality has a long history in cognitive science (Brunswik, 1955; Simon, 1990; Todd & Gigerenzer, 2012). The proposition that people are to some degree adapted to the problems they have to solve in the real world is maybe trivial. Figuring out how strong this adaptation is, on the other hand, has remained a big open question. Yet, it has been notoriously difficult to build models that are optimally adapted to the problems that people encounter in their everyday environment. From a technical perspective, the framework of meta-learning offers a solution to this problem. Yet, it has thus far only been applied to artificially-generated environments (Kumar et al., 2020; Binz & Schulz, 2022b; Binz et al., 2022a; Lake & Baroni, 2023; Jagadish et al., 2023). The main obstacle up to now was that the number of available real-world data sets was insufficient for meta-learning. We have shown that one can overcome this obstacle by prompting LLMs to generate a large collection of ecologically valid category learning tasks. We have then used meta-learning to obtain models that are optimally adapted to them, leading to a class of ecologically rational models that we call ERMI.
@@
While it is possible to come up with a mathematical formulation that allows generating features that match a given statistic like correlation or linearity, it is unclear how to design an algorithm that generates category-learning tasks that match all real-world statistics measured and unmeasured. This is exactly where the strength of pre-trained LLMs lies. If queried correctly, LLMs can act as a single, easy-to-access, rich – possibly infinite – source of real-world data.
10 Limitations
There are some limitations to generating ecological data that are worth noting. Firstly, it is uncertain whether the current approach can be scaled to generate several more orders of features and multi-class category learning problems. Secondly, strict data curation protocols are necessary if the approach is to be used in cases where unbiased factually correct data are required. Thirdly, preliminary analyses revealed that the data generated by LLMs are US/western-centric (Anthropic, 2023; Tamkin et al., 2023). Therefore, culturally diverse LLMs are needed to generate data that capture the rich diversity of the real world.
On the cognitive science front, there were also facets of human category learning not captured by ERMI. In particular, we found that ERMI generally learned faster than people. In the third experiment, for instance, ERMI already displayed the generalization patterns shown in Figure 4 after two blocks, while people required 16 blocks or more (Johansen & Palmeri, 2002). We believe that this gap can (at least partially) be closed by incorporating limited computational resources (Binz et al., 2022a; Jagadish et al., 2023) or other architectural constraints (Achterberg et al., 2023).
11 Conclusion
We have shown that LLMs can generate ecologically valid category learning tasks that can be used for meta-learning. With these models at hand, we then demonstrated that one can explain human category learning to a large extent using the principle of ecological rationality. Furthermore, the priors acquired by ERMI are rich enough that it achieves state-of-the-art performance on a real-world classification benchmark. In future work, we plan to scale up ERMI’s architecture, train it on classification tasks with a flexible number of features, increase the maximum number of data points, and allow for more than two classes.
Acknowledgements
We thank all reviewers for their constructive and thoughtful feedback. We would also like to thank the authors of Devraj et al. (2021), Nosofsky et al. (1994a), and Badham et al. (2017) for making the data from their study available. Furthermore, we thank the members of the “Computational Principles of Intelligence Laboratory” (CPI Lab), participants of the “Analytical Connectionism” summer school, Dan John, Yashas Annadani, and Laura Heidiri for their comments, discussions, and support. This work was supported by the Max Planck Society, the Volkswagen Foundation, and funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC2064/1–390727645.15/18.
Impact Statement
This paper presents work whose goal is to show that methods from machine learning, specifically large language models and meta-learning, can be used to advance our understanding of human cognition. An important finding from our work that might be of significance to a broader audience is that training on a large corpus of human-generated text has allowed large language models to capture certain key ecological features. Therefore, one has to be aware that such very powerful features can emerge in foundation models when scaled appropriately. There may be other potential societal consequences of our work, but we do not feel that they need to be specifically highlighted here.
@@ B.2 Data pre-processing, model architecture, and training
We filter out all tasks with more than two unique category labels and then binarize the category labels which are originally strings to make them consistent across tasks. The assignment of category labels, that is either ‘0’ or ‘1’, within a category learning task was randomized during batch creation. This ensures that there can be no unintended correlations between the stimuli seen during training and the labels (across all training data each input vector is assigned half of the time to label ‘0’ and half of the time to label ‘1’). We also normalized each feature independently using a min-max normalization scheme such that values taken by any feature lie always between zero and one. Both the task features and data points were shuffled while generating tasks. Note that the tasks generated by LLMs are typically of different lengths. Whenever the sampled tasks are of variable lengths, they are padded with zeros to match the length of the longest task sample within the batch. We additionally also sampled LLM-generated data points with replacement to match the length of the experimental task used in the Devraj et al. (2021) and Johansen & Palmeri (2002) studies. We resorted to this strategy as the LLM-generated tasks had a maximum of about 200 data points per task and by resampling, we can evaluate the model on experiments with larger horizons without any drop in performance. The batch size was set to 64 for three- and four-dimensional stimuli and to 32 for six-dimensional stimuli and it operated under a maximum steps regime of 400, 300, and 650 for three, four, and six-dimensional tasks respectively.
@@
The task features were mapped onto a 64-dimensional embedding space and positional encoded using sine and cosine functions of different frequencies as in Vaswani et al. (2017). A causal attention mask was then generated for the inputs such that the model makes conditional predictions on all preceding data points. The inputs along the attention mask are then passed to the transformer decoder model which had six layers, a model dimension of 64, 256 hidden units in the feed-forward network, and eight attention heads. The output of the transformer was then passed through a linear readout and sigmoid function to generate probability estimates for category 1. In practice, inference for all time steps is performed in parallel by passing a causal attention mask to the TransformerDecoder module in PyTorch (Paszke et al., 2019). We used binary cross-entropy (BCE) loss for a given batch of inputs and updated the model parameters using the ADAM optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-4}$. We trained all our models for a total of 500 000500000500\,000500 000 episodes.
@@ Appendix C LLMs generate ecologically valid category learning tasks
Features values:
We compared the distributions of input values generated by LLMs to input values from real-world datasets and found them to be extremely similar, as shown in Figure 7. We furthermore checked if there were spurious values in the LLM-generated data such as peaks at multiples of 5s and 10s, and found them to be at chance level (20% for multiples of 5s, 10% for multiples of 10s).
Additional analysis:
We furthermore computed the KL divergence between the histograms of the real-world data set and the synthetically generated data sets for each investigated metric. The resulting KL divergences confirm that the LLM-generated data provides a good proxy for real-world data (see Table 3). In Figure 8, we overlaid the histograms of the real-world data set and the synthetically generated data sets for the four different statistics discussed in the main paper.
Table 3: KL divergence between real-world data classification data and the synthetically generated data for different statistical measures
Real world tasks vs. Input Correlation Gini Coefficient Posterior Probability Features
LLM-generated $0.2716$ $0.2286$ $0.0202$ $0.0418$
MI Prior $7.7637$ $2.8170$ $2.5265$ $1.6758$
PFN Prior $7.7746$ $8.6391$ $2.5265$ $1.6775$
Figure 7: Comparing input features of real-world classification tasks and LLM-generated tasks: (a) Histogram of normalized input features from 28 different real-world binary classification tasks from the OpenML-CC18 benchmarking suite, (b) from ecologically valid category learning tasks generated from Claude-v2, (c) synthetic category learning tasks derived using the Bayesian logistic regression prior that were used to train the meta-learned inference (MI) model and (d) synthetic category learning tasks with nonlinear decision boundary derived using the Bayesian neural network prior that were used to train prior-fitted networks (PFN) model.
Figure 8: Comparing the ecological validity of different category learning tasks: Histogram of normalized input features (first column). Histogram of Pearson’s correlation coefficients computed between pairs of features (second column). Histogram of Gini coefficients computed over the logistic regression weights (third column). Linearity of the category learning task (fourth column). These four statistics were (a) ecologically valid category learning tasks generated from Claude-v2, (b) synthetic category learning tasks derived using the Bayesian logistic regression prior that were used to train the meta-learned inference (MI) model, and (c) synthetic category learning tasks with nonlinear decision boundary derived using the Bayesian neural network prior that were used to train prior-fitted networks (PFN) model were overlaid on the same statistics computed on 28 different real-world binary classification tasks from the OpenML-CC18 benchmarking suite.
@@ Appendix D Cognitive models
Claude-v2:
We used Claude-v2 as a cognitive model of human category learning. To do that, we prompted Claude-v2 by paraphrasing the instructions given to participants, see Appendix G and H for task-specific prompts. In essence, the model was instructed via the prompt that it would be presented with stimuli belonging to two categories and would receive feedback regarding the correct category following each stimulus. It was instructed to learn a rule based on stimulus features to assign the stimulus to the correct category with increasing experience. As the Claude API only returns the Claude-v2 output token and not its probability, the model’s prediction was coded as a binary variable, $\pi(k=1\mid x_{t};x_{1:t-1},y_{1:t-1})$. The final model predicted category probabilities is again a mixture between the category prediction from the model and a random guess as mentioned in Equation 9.
$p(k=1\mid x_{t})=(1-\epsilon)\pi(k=1\mid x_{t};x_{1:t-1},y_{1:t-1})+\epsilon\cdot\mathrm{K}^{-1}$ (9)
@@ Appendix E Fitting models to human data
Claude-v2:
We used the same procedure as above except that only the guessing parameter, $\epsilon$, is learned.
@@ G.3 Additional observations and results
Table 5: Mean performance of humans and models for each rule type in replication of Shepard et al. (1961) study over 15 blocks. Human data was taken from Table 1 in Nosofsky et al. (1994a). Details of model simulations can be found in Appendix G.
Model Rule MSE
Type 1 Type 2 Type 3 Type 4 Type 5 Type 6
Humans .0201 .0565 .1015 .1120 .1212 .2048 .0000
ERMI .0586 .0891 .0855 .0826 .0888 .1172 .0287
MI .0686 .4089 .2404 .1431 .2880 .4201 .2627
PFN .0170 .3405 .1533 .0226 .2371 .3975 .1736
RMC .1329 .2215 .1903 .1718 .2132 .3364 .1003
G.4 Claude-v2 as a cognitive model of human category learning
Simulations:
To run simulations of the Badham et al. (2017) study on Claude-v2, we queried it using the prompt shown in Appendix G.4. The geometric stimuli used in the experiment are described in textual format. The order of stimulus presentation was randomized, and the number of stimulus presentations per block was matched to the original study. We ran 96 simulation runs for each of the six rules.
[…]
Figure 10: Unlike ERMI, Claude-v2 does not show human-like learning difficulties: (a-c) Average error probabilities for each task type in each block of 16 trials for (a) humans, (b) ERMI, and (c) LLM. Human data in (a) was reproduced from Table 1 in Nosofsky et al. (1994a). ERMI was simulated on type 1-6 tasks for 50 runs with the inverse temperature set to $\beta=0.4$. Claude-v2 was simulated for 94 runs each on type 1-6 tasks with temperature term set to 0.
Table 6: Mean performance of humans and models for each rule type in the Badham et al. (2017) study. Details of model simulations can be found in Appendix G.
Model Rule MSE
Type 1 Type 2 Type 3 Type 4 Type 5 Type 6
Human .0460 .1267 .2157 .2307 .2297 .3003 .0000
ERMI .0912 .1374 .1317 .1248 .1361 .1798 .0538
LLM .0206 .0521 .0404 .0368 .0482 .0752 .1771
MI .0959 .4337 .2655 .1657 .3175 .4387 .1795
PFN .0402 .3908 .1887 .0527 .2722 .4237 .1541
RMC .1593 .2744 .2314 .2164 .2576 .3845 .0564
@@ H.3 Additional observations and results
A better fit of the prototype model to the MI and PFN in the latter half of the experiment (as shown in Figure 3 and 11) suggests that, as observed in Shepard’s task, they are not able to learn exceptions like the prototype model as mentioned before.
H.4 Claude-v2 as a cognitive model of human category learning
Simulations:
To run simulations of the Smith & Minda (1998) study on Claude-v2, we queried it using the prompt shown in Appendix H.4. The nonsense stimuli used in the experiment are provided in text. The order of presentation of the stimulus was also randomized, and the number of stimulus presentations in a block was matched to the original study. We ran the model for a total of 120 simulation runs.
[…]
Figure 12: Claude-v2 performs purely exemplar-based category learning: (a-c) The average error of exemplar- and prototype-based models fitted to (a) human choices, (b) simulated choices from ERMI, and (c) simulated choices from Claude-v2 for each block of 56 trials. Human data in (a) was reproduced from Smith & Minda (1998). ERMI was simulated using inverse temperature values fitted to participants’ choices in Devraj et al. (2021). The mean of the fitted inverse temperature and its standard error for ERMI was $0.09\pm 0.01$. Claude-v2 was queried using the prompt shown in Appendix H.4 with the temperature term set to 0 for 120 runs. The shaded region shows the standard error of the mean.